a fish was caught whos tail weighs 9 pounds. His head weighed as much as his tail and half his body,and his body weighs as much as his head and his tail. how much does the fish weigh

Answer by Abbey(339)

let's break this apart into manageable pieces:
a fish was caught whos tail weighs 9 pounds.
Let the Tail (T) = 9 pounds
His head weighed as much as his tail and half his body,
Head (H) = T + 1/2 Body (B)
H=9+B/2
and his body weighs as much as his head and his tail.
B=H+T
B=H+9

how much does the fish weigh
since we know that the head =9+b/2, we can substitute into our second equation:
B=9+b/2+9
B=18+b/2
b-b/2=18
b/2=18
b=36
so the body weighs 36 pounds,
the head weighs 27 pounds
and the tail weighs 9 pounds
the fish (in total) weighed 72 pounds
That is certainly a whopper!